OpenStudy (anonymous):

@neer2890 can you solve this? 2/3 (x+6) ..

OpenStudy (neer2890):

\[\frac{ 2 }{ 3 }(x+6)\] =\[\frac{ 2 }{ 3 }x+\frac{ 2 }{ 3 }*6=\frac{ 2 }{ 3 }x+4\]

OpenStudy (anonymous):

quadratic?

Miracrown (miracrown):

this is just an expression with 2/3 multiplying (x+6).

Miracrown (miracrown):

Is there something more to do with this expression? @gua

OpenStudy (anonymous):

give the product...

Miracrown (miracrown):

- the only thing we can do here is distribute the 2/3 inside the ( ):  Multiplying the 2/3 times 6 gives us the 4.